r A badminton birdie travels at 25m/s towards a person. The person swats the birdie and it leaves the racquet at 15 m/s. If the birdie's change in velocity took only 0.5 seconds what is the acceleration of the birdie while being hit.
1
Expert's answer
2017-10-13T15:06:06-0400
Before the impact, let the velocity of the baseball was v=25 m/s.
After being hit by the bat its velocity is -u=-15 m/s
(the minus sign indicates that the direction was reversed).

So, change in velocity,
∆v=v-(-u)=25+15=35 m/s
Acceleration is defined as the rate of change in velocity, i.e. actual change in velocity divided by the time taken to change it. Time taken to change velocity is the time of actual contact of the racquet and birdie, i.e. 0.5 s.
a=∆v/∆t=35/0.5=70 m/s^2